jigglypuff314 (jigglypuff314):

A calc problem... If f(x) = 3secx - 2cosx + 4, evaluate the limit as x->(4pi/3) (f(x)-f(4pi/3))/(x-(4pi/3))

jigglypuff314 (jigglypuff314):

If f(x) = 3secx - 2cosx + 4, evaluate the limit \[\lim_{x \rightarrow 4\pi /3} \frac{ f(x)-f(4\pi /3) }{ x - (4\pi /3) }\]

OpenStudy (anonymous):

You can do it. It's a derivative.

jigglypuff314 (jigglypuff314):

is it -5√3 ?

OpenStudy (anonymous):

not even sure how you got that. Where are you stuck anyway?

jigglypuff314 (jigglypuff314):

why don't you try plugging in and look for the limit first this was a bonus question on my test... I think my teacher was trolling...

jigglypuff314 (jigglypuff314):

I got the answer that I got by finding the derivative then plugging in (4pi/3) but not sure how to get the answer by limit process 0.o

OpenStudy (anonymous):

Well, it probably takes a lot of algebra that sucks to do.